Contribution Margin, Break-Even Units, Contribution Margin Income Statement, Margin of Safety

Zebra Company manufactures custom-designed skins (covers) for iPods® and other portable MP3 devices. Variable costs are $13.00 per custom skin, the price is $20, and fixed costs are $205,800.

Required:

1. What is the contribution margin for one custom skin? Round your answer to the nearest cent.
$ per custom skin

2. How many custom skins must Zebra Company sell to break even?
    custom skins

3. If Zebra Company sells 30,000 custom skins, what is the operating income?
$

4. Calculate the margin of safety in units and in sales revenue if 30,000 custom skins are sold.

Margin of safety in units units
Margin of safety in sales revenue $
Answer #1

1. Contribution margin = Selling price - variable cost = $ 20 - $ 13 = $ 7.

2. Break even units = Fixed cost / CM per unit = $ 205800 / $ 7 = 29400 units.

3. Operating income = ( 30000 * 7 - 205800 ) = $ 4200.

4. Margin of safety units = Sales units - Break even units = 30000 - 29400 = 600 units.

>> Margin of safety sales = Margin of safety units * selling price = 600 * $ 20 = $ 12000.